# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

### v1

@@ 2.3. Semiclassical definitions
Following [DZ19]*§E.1.7 and [Zw12]*§14.2.2, we can generalize this quantization to pseudodifferential operators on manifolds $\Psi^{k}_{h}(M)$. We denote the principle symbol map by
@@
Semiclassical measures are geodesic-flow invariant probability measure with support contained in $S^{*}M$ (see [Zw12]*§§5.2-1).
@@ 2.5. Symbol class
Recall the definitions of the Lagrangian foliations $L_{s}$ and $L_{u}$ from (2.12). From the argument in [DJ18]*§2.3, if $a\in C_{c}^{\infty}(T^{*}M\setminus 0)$ is an $h$-independent symbol,
(2.19) $a\circ\varphi_{t}\in S^{\operatorname{comp}}_{L_{s},\rho,\rho^{\prime}}(T^{*}M\setminus 0)\text{ and }a\circ\varphi_{-t}\in S^{\operatorname{comp}}_{L_{u},\rho,\rho^{\prime}}(T^{*}M\setminus 0)\text{ uniformly in }t\in[0,\rho\log h^{-1}],$
for any $\rho^{\prime}$ such that $0\leq\rho^{\prime}\leq\tfrac{\rho}{2}$ and $\rho+\rho^{\prime}<1$. This follows from using (2.11) to rewrite the derivative bounds (2.18) in terms of the frame in (2.6), then using the commutation relations in (2.7).
@@
We remark that the full version of  [DJ18]*Lemma A.8 holds only for hyperbolic surfaces, however the proof for the part quoted above holds in all dimensions.
@@ 3.2. Group theoretic preliminaries
A choice of minimal parabolic $P^{\pm}$ which has unipotent radical $N^{\pm}$ is therefore given by $P^{\pm}=K_{0}AN^{\pm}$, with $K_{0}$ as in (2.5); all other minimal parabolics are $G$-conjugate to $P^{\pm}$. This decomposition of $P^{\pm}$ is known as the Langland’s decomposition, where in the literature typically one uses “$M$” as opposed to “$K_{0}$”. Importantly, $N_{G}(N^{\pm})=P^{\pm}$ and there is an isomorphism of $N^{\pm}$ with $\mathbb{R}^{n}$, induced by sending an element $b\in N^{\pm}$ to the corresponding vector $\vec{v}$, which intertwines the action by conjugation of $K_{0}$ on $N^{\pm}$ and the linear action of $K_{0}\cong\operatorname{SO}(n)$ on $\mathbb{R}^{n}$. Moreover, under such an isomorphism, the conjugation action by $A$ on $N^{\pm}$ is intertwined with the action of $\mathbb{R}$ on $\mathbb{R}^{n}$ by scaling; precisely, by scaling by a factor of $\cosh(t)+\sinh(t)$.
We use these facts to deduce a few group theoretic lemmas, which we require in the sequel. All of these facts are well-known, but we provide proofs for completeness.
@@ 3.3. Classifying $U^{\pm}$ orbit closures in $F^{*}M$
Note that $AU^{\pm}$ is not generated by unipotents so we cannot apply Ratner’s theorem to classify its orbits as in Lemmas 3.6 and 3.7. Therefore the proof of Proposition 3.8 will require additional understanding of the failure of equidistribution of $U^{\pm}$-orbits inside $W_{2}$-orbits, which is classified in the work of Ratner [RatnerMeasure, RatnerTop] and Dani–Margulis [DM]. We opt to give a complete account of the relevant details herein and note the similarities between some arguments found in [LeeOh], [BFMS].
@@ 4. Fractal uncertainty principle
We spend the rest of this section generalizing Theorem  4.3. We do so by adapting the work of  [BD18]*§§2.2, 4 to higher dimensions. Notably, Han and Schlag in  [HS20] adapted the work of  [BD18] to higher dimensions, proving a higher-dimensional fractal uncertainty principle, then generalizing said fractal uncertainty principle similarly to how we will generalize Theorem  4.3. However, their higher-dimensional fractal uncertainty principle requires stronger conditions than Theorem  4.3: while $X_{-}$ must satisfy a condition equivalent to ball porosity, $X_{+}$ must satisfy a stronger condition than line porosity. Namely, $X_{+}$ must be covered by a finite number of sets $Y_{1},\ldots,Y_{m}$ where each $Y_{i}$ is the product of $n$ 1-dimensional $\delta$-regular sets. Thus, although  [HS20] has many similarities with the work in this section, due to the different conditions on $X_{+}$, we cannot use their results.
@@ 4.1. Porosity properties
The first step in generalizing Theorem  4.3 is to adapt  [BD18]*Lemmas 2.1-4 to higher dimensions to show that line and ball porosity are preserved under certain operations (although the porosity constant $\nu$ may decrease and the scales may change). Notably,  [BD18] used $\delta$-regularity instead of porosity and thus used different proofs to ours.
First we show that porosity is preserved under affine transformations.
@@
Let $X$ be $\nu$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$. Fix $\lambda>0$ and $y\in\mathbb{R}$. Then $\tilde{X}\coloneqq y+\lambda X$ is $\nu$-porous on balls from scales $\lambda\alpha_{0}$ to $\lambda\alpha_{1}$.
@@
Let $X$ be $\nu$-porous on lines from scales $\alpha_{0}$ to $\alpha_{1}$. Fix $\lambda>0$ and $y\in\mathbb{R}$. Then $\tilde{X}\coloneqq y+\lambda X$ is $\nu$-porous on lines from scales $\lambda\alpha_{0}$ to $\lambda\alpha_{1}$.
@@
$C_{1}^{-1}|x-y|\leq|\varkappa(x)-\varkappa(y)|\leq C_{1}|x-y|.$
Let $X\subset\mathbb{R}^{n}$ and suppose $\varkappa(X)$ is $\nu$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$. Then $X$ is $\frac{\nu}{C_{1}^{2}}$-porous on balls from scales $C_{1}\alpha_{0}$ to $C_{1}\alpha_{1}$.
@@
Let $x\in\mathbb{R}^{n}$ and pick $C_{1}\alpha_{0}<R<C_{1}\alpha_{1}$. It suffices to show that there exists $y\in B_{R/2}(x)$ such that $B_{\nu R/C_{1}^{2}}(y)\cap X=\emptyset$.
@@
$B_{\frac{R}{2C_{1}}}(\varkappa(x))\subset\varkappa\left(B_{\frac{R}{2}}(x)\right).$
By the ball porosity of $\varkappa(X)$, there exists $\varkappa(y)\in B_{R/2C_{1}}(\varkappa(x))$ such that $B_{\nu R/C_{1}}(\varkappa(y))\cap\varkappa(X)=\emptyset.$ Thus,
$\varkappa^{-1}\left(B_{\frac{\nu R}{C_{1}}}(\varkappa(y))\right)\cap X=\emptyset.$
Finally, since $B_{\nu R/C_{1}^{2}}(y)\subset\varkappa^{-1}(B_{\nu R/C_{1}}(\varkappa(y)))$, we know $B_{\nu R/C_{1}^{2}}(y)\cap X=\emptyset$. The proof concludes by noting that $y\in B_{R/2}(x)$. ∎
@@ 4.2. Generalizations of the fractal uncertainty principle
We now follow the argument of  [BD18]*§§4.1-2, generalizing Lemma  4.11 first to operators with variable amplitude, then to operators with a general phase.
@@
We see that $B$ is a Fourier integral operator associated to $\kappa$.
@@
Choose $y^{\prime}\in\mathbb{R}^{n}$ such that for some $x^{\prime}\in\mathbb{R}^{n}$, $(x^{\prime},y^{\prime})\in U$. By  (4.7), there exists some neighborhood $U_{x^{\prime}}$ of $x^{\prime}$ with $U_{x^{\prime}}\times\{y^{\prime}\}\subset U$ where $\partial_{y}\Phi(\cdot,y^{\prime})$ is locally invertible. Therefore, for any $x_{1},x_{2}\in U_{x^{\prime}}$,
$\frac{|\partial_{y}\Phi(x_{1},y^{\prime})-\partial_{y}\Phi(x_{2},y^{\prime})|}{|x_{1}-x_{2}|}\leq\sup_{x\in U_{x^{\prime}}}\|\partial^{2}_{xy}\Phi(x,y^{\prime})\|,$
@@
$\frac{|x_{1}-x_{2}|}{|\partial_{y}\Phi(x_{1},y^{\prime})-\partial_{y}\Phi(x_{2},y^{\prime})|}\leq\sup_{x\in U_{x^{\prime}}}\|\partial_{x}(\partial_{y}\Phi(x,y^{\prime}))^{-1}\|=\sup_{x\in[\partial_{y}\Phi(U_{x^{\prime}},y^{\prime})]^{-1}}\|(\partial_{xy}^{2}\Phi(x,y^{\prime}))^{-1}\|.$
From  (4.7), we see $\sup_{x\in U_{x^{\prime}}}\|\partial^{2}_{xy}\Phi(x,y^{\prime})\|$ and $\sup_{x\in[\partial_{y}\Phi(U_{x^{\prime}},y^{\prime})]^{-1}}\|(\partial_{xy}^{2}\Phi(x,y^{\prime}))^{-1}\|$ are nonzero and depend continuously on $x^{\prime}$ and $y^{\prime}$. Therefore, there exists $C=C_{x^{\prime},y^{\prime}}>0$ and a neighborhood $U_{y^{\prime}}$ of $y^{\prime}$ with $U_{x^{\prime}}\times U_{y^{\prime}}\subset U$ such that for $y\in U_{y^{\prime}}$,
@@
Thus, breaking the symbol $b$ into pieces using a partition of unity, we may assume that
@@
where $D_{-},D_{+},D^{\prime}_{+}$ are balls with $D_{+}^{\prime}\Subset D_{+}$ and for $(x_{1},y),(x_{2},y)\in D_{-}\times D^{\prime}_{+}$,  (4.8) holds. We assume that $J\subset D_{+}$, else for $h$ sufficiently small, $\mathds{1}_{X_{-}(h^{\rho/2})}B(h)\mathds{1}_{X_{+}(h^{\rho})\cap J}=0$.
Let $y_{0}$ be the center of $J$ and define the function
@@
Using (4.8), from Lemma  4.9, we see $\tilde{X}_{-}$ is $\frac{\nu}{C^{2}}$-porous on balls from scales $Ch^{\varrho}$ to $1$. Thus, $\tilde{X}_{-}$ is $\frac{\nu}{C^{2}}$-porous on balls from scales $\tilde{h}^{2\varrho-1}$ to $1$. From Lemma  4.5, $\tilde{X}_{+}$ is $\nu$-porous on lines from scales $\tilde{h}^{2\varrho-1}$ to $1$.
@@ 5.1. Partition of unity
We now construct $A_{1},A_{2}$. From Lemma  2.6, there exists $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$ such that $\chi_{1}+\chi_{2}=1$, $\operatorname{supp}\chi_{1}\subset U_{\mu}$, and $S^{*}M\setminus\operatorname{supp}\chi_{1}$, $S^{*}M\setminus\operatorname{supp}\chi_{2}$ are $U_{1}^{-}$-dense. By Lemma 2.4, there exists compact $U_{1}^{-}$-dense sets $U_{1},U_{2}\subset S^{*}M$ such that $U_{1}\Subset S^{*}M\setminus\operatorname{supp}\chi_{1}$ and $U_{2}\Subset S^{*}M\setminus\operatorname{supp}\chi_{2}$. We extend $\chi_{1}$ and $\chi_{2}$ to be homogeneous functions of order 00 on $T^{*}M\setminus 0$.
@@ 5.2. Dynamical refinement of partition of unity
We outline the proof of Lemma 5.2 to provide intuition, but defer the full proof to  [DJ18]. The proof in [DJ18] uses [DJ18]*Lemma A.8, which is a statement about hyperbolic surfaces. However, the only part of [DJ18]*Lemma A.8 used is (2.20), which holds in all dimensions.
Recall that uniformly in $t\in[0,\rho\log h^{-1}]$, from (2.19), we know $a\circ\varphi_{t}\in S^{\operatorname{comp}}_{L_{s},\rho,0}(T^{*}M\setminus 0)$ and from (2.20), we have an Egorov’s theorem. Thus, we can use the following result, which combines the statements of [DJ18]*Lemma A.1 and [DJ18]*Lemma A.6.
@@
We adapt  [DJ18]*§4, which in turn used many of the tools from  [An08]*§2.

### v2

@@ 2.3. Semiclassical definitions
Following [DZ19]*§E.1.7 and [Zw12]*§14.2.2, we can generalize this quantization to pseudodifferential operators on manifolds $\Psi^{k}_{h}(M)$. We denote the principal symbol map by
@@
Semiclassical measures are geodesic-flow invariant probability measures with support contained in $S^{*}M$ (see [Zw12]*§§5.2–1).
@@ 2.5. Symbol class
Recall the definitions of the Lagrangian foliations $L_{s}$ and $L_{u}$ from (2.12). From the argument in [DJ18]*§2.3, if $a\in C_{c}^{\infty}(T^{*}M\setminus 0)$ is an $h$-independent symbol, uniformly in $t\in[0,\rho\log h^{-1}]$,
(2.19) $a\circ\varphi_{t}\in S^{\operatorname{comp}}_{L_{s},\rho,0}(T^{*}M\setminus 0)\text{ and }a\circ\varphi_{-t}\in S^{\operatorname{comp}}_{L_{u},\rho,0}(T^{*}M\setminus 0).$
This follows from using (2.10) to rewrite the derivative bounds (2.18) in terms of the frame in (2.6), then using the commutation relations in (2.7).
@@
We remark that [DJ18]*Lemma A.8 is stated for hyperbolic surfaces, however the proof holds in all dimensions.
@@ 3.2. Group theoretic preliminaries
A choice of minimal parabolic $P^{\pm}$ which has unipotent radical $N^{\pm}$ is therefore given by $P^{\pm}=K_{0}AN^{\pm}$, with $K_{0}$ as in (2.5); all other minimal parabolics are $G$-conjugate to $P^{\pm}$. This decomposition of $P^{\pm}$ is known as the Langland’s decomposition, where in the literature typically one uses “$M$” as opposed to “$K_{0}$”. Importantly, $N_{G}(N^{\pm})=P^{\pm}$ and there is an isomorphism of $N^{\pm}$ with $\mathbb{R}^{n}$, induced by sending an element $b\in N^{\pm}$ to the corresponding vector $\vec{v}$, which intertwines the action by conjugation of $K_{0}$ on $N^{\pm}$ and the linear action of $K_{0}\cong\operatorname{SO}(n)$ on $\mathbb{R}^{n}$. Moreover, under such an isomorphism, the conjugation action by $A$ on $N^{\pm}$ is intertwined with the action of $\mathbb{R}$ on $\mathbb{R}^{n}$ by scaling; precisely, by scaling by a factor of $\cosh(t)+\sinh(t)=e^{t}$.
We use these facts to deduce a few group theoretic lemmas, which we require in the sequel. All of these facts are standard, but we provide proofs for completeness.
@@ 3.3. Classifying $U^{\pm}$ orbit closures in $F^{*}M$
Note that $AU^{\pm}$ is not generated by unipotents so we cannot apply Ratner’s theorem to classify its orbits as in Lemmas 3.6 and 3.7. Therefore the proof of Proposition 3.8 will require additional understanding of the failure of equidistribution of $U^{\pm}$-orbits inside $W_{2}$-orbits, which is classified in the work of Ratner [RatnerMeasure, RatnerTop] and Dani–Margulis [DM]. Many of the arguments below are inspired by [LeeOh], which develops a topological approach to the study of orbit closures in $\operatorname{SO}_{0}(1,n+1)$; there are also some similarities with arguments found in [BFMS]. For completeness, we opt to give a complete account of the relevant details in what follows.
@@ 4. Fractal uncertainty principle
We spend the rest of this section generalizing Theorem  4.3. We do so by adapting the work of  [BD18]*§§2.2, 4 to higher dimensions.
@@ 4.1. Porosity properties
The first step in generalizing Theorem  4.3 is to adapt  [BD18]*Lemmas 2.1–4 to higher dimensions to show that line and ball porosity are preserved under certain operations (although the porosity constant $\nu$ may decrease and the scales may change). Notably,  [BD18] used $\delta$-regularity instead of porosity and thus used different proofs to ours.
First we show that porosity is preserved under dilations and shifts.
@@
Let $X$ be $\nu$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$. Fix $\lambda>0$ and $y\in\mathbb{R}^{n}$. Then $\tilde{X}\coloneqq y+\lambda X$ is $\nu$-porous on balls from scales $\lambda\alpha_{0}$ to $\lambda\alpha_{1}$.
@@
Let $X$ be $\nu$-porous on lines from scales $\alpha_{0}$ to $\alpha_{1}$. Fix $\lambda>0$ and $y\in\mathbb{R}^{n}$. Then $\tilde{X}\coloneqq y+\lambda X$ is $\nu$-porous on lines from scales $\lambda\alpha_{0}$ to $\lambda\alpha_{1}$.
@@
Let $U\subset\mathbb{R}^{n}$ be a ball and $\varkappa:U\rightarrow\varkappa(U)$ be a diffeomorphism with
$C_{1}^{-1}|x-y|\leq|\varkappa(x)-\varkappa(y)|\leq C_{1}|x-y|,$
for $x,y\in U$. Let $X\subset U$ and suppose $\varkappa(X)$ is $\nu$-porous on balls from scales $\alpha_{0}$ to $\alpha_{1}$. Then $X$ is $\frac{\nu}{2C_{1}^{2}}$-porous on balls from scales $C_{1}\alpha_{0}$ to $C_{1}\alpha_{1}$.
@@
Let $x\in\mathbb{R}^{n}$ and pick $C_{1}\alpha_{0}<R<C_{1}\alpha_{1}$. We need to show that there exists $y\in B_{R/2}(x)$ such that $B_{\nu R/C_{1}^{2}}(y)\cap X=\emptyset$. Note that exists $\tilde{x}\in B_{R/2}(x)$ such that either $B_{R/4}(\tilde{x})\subset U\cap B_{R/2}(x)$ or $B_{R/4}(\tilde{x})\cap U=\emptyset$. Therefore, it suffices to assume that $B_{R/4}(x)\subset U$ and show there exists $y\in B_{R/4}(x)$ such that $B_{\nu R/2C_{1}^{2}}(y)\cap X=\emptyset$.
@@
$B_{\frac{R}{4C_{1}}}(\varkappa(x))\subset\varkappa\left(B_{\frac{R}{4}}(x)\right).$
By the ball porosity of $\varkappa(X)$, there exists $\varkappa(y)\in B_{R/4C_{1}}(\varkappa(x))$ such that $B_{\nu R/2C_{1}}(\varkappa(y))\cap\varkappa(X)=\emptyset.$ Thus,
$\varkappa^{-1}\left(B_{\frac{\nu R}{2C_{1}}}(\varkappa(y))\cap\varkappa(U)\right)\cap X=\emptyset.$
Finally, since $B_{\nu R/2C_{1}^{2}}(y)\subset\varkappa^{-1}(B_{\nu R/2C_{1}}(\varkappa(y))\cap\varkappa(U))\cup(\mathbb{R}^{n}\setminus U)$, $B_{\nu R/2C_{1}^{2}}(y)\cap X=\emptyset$. The proof concludes by noting that $y\in B_{R/4}(x)$. ∎
@@ 4.2. Generalizations of the fractal uncertainty principle
We now follow the argument of  [BD18]*§§4.1–2, generalizing Lemma  4.11 first to operators with variable amplitude, then to operators with a general phase.
@@
We see that $B$ is a semiclassical Fourier integral operator associated to $\kappa$.
@@
Breaking the symbol $b$ into pieces using a partition of unity, we may assume that $\operatorname{supp}b\subset D_{-}\times D_{+}\subset U$, where $D_{-},D_{+}$ are balls . We assume that $J\subset D_{+}$, else for $h$ sufficiently small, $\mathds{1}_{X_{-}(h^{\rho/2})}B(h)\mathds{1}_{X_{+}(h^{\rho})\cap J}=0$.
Let $y_{0}$ be the center of $J$. Choose $x_{0}\in D_{-}$. By  (4.7), there exists some neighborhood $U_{x_{0}}$ of $x_{0}$ with $U_{x_{0}}\subset D_{-}$ where $\partial_{y}\Phi(\cdot,y_{0})$ is locally invertible. Therefore, for any $x_{1},x_{2}\in U_{x_{0}}$,
$\frac{|\partial_{y}\Phi(x_{1},y_{0})-\partial_{y}\Phi(x_{2},y_{0})|}{|x_{1}-x_{2}|}\leq\sup_{x\in U_{x_{0}}}\|\partial^{2}_{xy}\Phi(x,y_{0})\|,$
@@
$\frac{|x_{1}-x_{2}|}{|\partial_{y}\Phi(x_{1},y_{0})-\partial_{y}\Phi(x_{2},y_{0})|}\leq\sup_{x\in U_{x_{0}}}\|\partial_{x}(\partial_{y}\Phi(x,y_{0}))^{-1}\|=\sup_{x\in[\partial_{y}\Phi(U_{x_{0}},y_{0})]^{-1}}\|(\partial_{xy}^{2}\Phi(x,y_{0}))^{-1}\|.$
From  (4.7), we see $\sup_{x\in U_{x_{0}}}\|\partial^{2}_{xy}\Phi(x,y_{0})\|$ and $\sup_{x\in[\partial_{y}\Phi(U_{x_{0}},y_{0})]^{-1}}\|(\partial_{xy}^{2}\Phi(x,y_{0}))^{-1}\|$ are nonzero and depend continuously on $x_{0}$ and $y_{0}$. Therefore, there exists $C=C_{x_{0},y_{0}}>0$ and a neighborhood $U_{y_{0}}$ of $y_{0}$ with $U_{x_{0}}\times U_{y_{0}}\subset U$ such that for $y\in U_{y_{0}}$,
@@
Thus, using another partition of unity to further decrease the support of $b$ and shrinking $D_{-}$, we may assume that
@@
where $D^{\prime}_{+}$ is a ball with $D_{+}^{\prime}\Subset D_{+}$ and for $(x_{1},y),(x_{2},y)\in D_{-}\times D^{\prime}_{+}$,  (4.8) holds.
Now define the function
@@
Using (4.8), from Lemma  4.9, we see $\tilde{X}_{-}$ is $\frac{\nu}{2C^{2}}$-porous on balls from scales $Ch^{\varrho}$ to $1$. Thus, $\tilde{X}_{-}$ is $\frac{\nu}{2C^{2}}$-porous on balls from scales $\tilde{h}^{2\varrho-1}$ to $1$. From Lemma  4.5, $\tilde{X}_{+}$ is $\nu$-porous on lines from scales $\tilde{h}^{2\varrho-1}$ to $1$.
@@ 5.1. Partition of unity
We now construct $A_{1},A_{2}$. From Lemma  2.7, there exists $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$ such that $\chi_{1}+\chi_{2}=1$, $\operatorname{supp}\chi_{1}\subset U_{\mu}$, and $S^{*}M\setminus\operatorname{supp}\chi_{1}$, $S^{*}M\setminus\operatorname{supp}\chi_{2}$ are $U_{1}^{-}$-dense. By Lemma 2.4, there exists compact $U_{1}^{-}$-dense sets $U_{1},U_{2}\subset S^{*}M$ such that $U_{1}\Subset S^{*}M\setminus\operatorname{supp}\chi_{1}$ and $U_{2}\Subset S^{*}M\setminus\operatorname{supp}\chi_{2}$. We extend $\chi_{1}$ and $\chi_{2}$ to be homogeneous functions of order 00 on $T^{*}M\setminus 0$.
@@ 5.2. Dynamical refinement of partition of unity
We outline the proof of Lemma 5.2 to provide intuition, but defer the full proof to  [DJ18]. Recall that uniformly in $t\in[0,\rho\log h^{-1}]$, from (2.19), we know $a\circ\varphi_{t}\in S^{\operatorname{comp}}_{L_{s},\rho,0}(T^{*}M\setminus 0)$ and from (2.20), we have an Egorov’s theorem. Thus, we can use the following result, which combines the statements of [DJ18]*Lemma A.1 and [DJ18]*Lemma A.6.
@@
We adapt  [DJ18]*§4, which in turn used many of the tools from  [An08]*§2. The following argument could be easily modified to hold for $o(h/\log h^{-1})$ quasimodes.